OpenStudy (anonymous):

7n – (4 – 2n) = 3(n + 5) - 1 solve:

Parth (parthkohli):

First distribute: \( \color{Black}{\Rightarrow 7n - 4 + 2n = 3n + 15 }\) \( \color{Black}{\Rightarrow 11n - 4 = 3n + 15 }\) Can you do it now?

Parth (parthkohli):

I meant \(9n - 4 = 3n + 15\)

OpenStudy (anonymous):

add 4 right?

Parth (parthkohli):

yes :)

Parth (parthkohli):

Then subtract 6n and divide both sides by 6.

sam (.sam.):

\[\begin{array}{l} 9 n-4=3 (n+5)-1 \\ \text{Expand out terms on the right hand side:} \\ 9 n-4=3 n+14 \\ \text{Add }(4-3 n)\text{ to both sides:} \\ 6 n=18 \\ \text{Divide both sides by 6 and that's all } \\\end{array}\]

OpenStudy (anonymous):

so 3 thank you !

Parth (parthkohli):

lol I missed out that -1

OpenStudy (anonymous):

woops lol